OpenStudy (anonymous):

Solve the following system of equations. x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1

OpenStudy (anonymous):

x=-3,y=-1,z=-5

OpenStudy (chaise):

I'm really bad at these - but I will do my best, for you. Sorry if I stuff up. x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 Make x the subject in the first equation. 3y-2z-16=x Substitute this value into your second and third equation. 2(3y-2z-16) – y + 2z = –15 2(3y-2z-16) – 2y – z = 1 Distribute. 2(3y-2z-16) – y + 2z = –15 2(3y-2z-16) – 2y – z = 1 6y-4z-32-y+2z=-15 6y-4z-32-y-z=1 Gather like terms 5y-2z-32=-15 5y-5z-32=1 Plus 32 both sides. 6y-2z=17 <-- multiply by 5 5y-5z=31 <-- multiply by 2 30y-10z=17 10y-10z=62 Take away to eliminate: 20y=-45 y = -9/4 Sub this back in. 10y-10z=62 10x(-9/4)-10z=62 -22.5-10z=62 -10z = 84.5 z = -8.45 2x – 2(-9/4) – (-8.45) = 1 2x+79/20 =1 2x =-59/20 x = -2.95 x = -2.95 y = -9/4 z - 8.45 Sorry if this is wrong. I tried ):

OpenStudy (chaise):

This is wrong - I will leave it here though, so everyone can glory over my failure, or find where I went wrong. Sorry ):

OpenStudy (anonymous):

are you allowed to use matrices to solve this Psssst? It would be easier with matrices.

OpenStudy (anonymous):

Not yet. I use that in the next lesson.

OpenStudy (anonymous):

Make x the subject in the first equation. 3y-2z-16=x I think here is the error or am I wrong? x=-16-3y-2z

OpenStudy (chaise):

More than likely - I think the procedure is correct though, I hope. Sorry again :x

OpenStudy (anonymous):

x + 3y + 2z = –16 [1] 2x – y + 2z = –15 [2] 2x – 2y – z = 1 [3] [2] - [3] : y + 3z = -16 [4] [1]x2: 2x + 6y + 4z = -32 [5] [2] - [5]: -7y -2z = 17 [6] we now solve [4] with [6]: [4]x7: 7y + 21z = -112 [7] [5] + [7]: 19z = -95 hence z = -5. From [4], y -15 = -16, hence y = -1. From [1], x - 3 - 10 = -16 hence x = -3

OpenStudy (anonymous):

xactxx that's fine

OpenStudy (anonymous):

you can also solve these by method of elimination subtract: 1 - 2: -x + 4y = -1.............................(a) multiply 3rd equn. by 2 4x-4y-2z = 2 add this to 2nd equn. 6x -5y = -13..............................(b) x (a) by 6: -6x + 24y = -6 add this to (b): 19y = -19 y = -1 substitute y = -1 into (a) -x -4 = -1 x = -3 find z by plugging y and x in first equn.: -3 +3*-1 + 2z = -16 2z = -16 + 6 = -10 z = -5 x=-3,y=-1,z=-5